# Stagewise Reinforcement Learning and the Geometry of the Regret Landscape

### v1

@@ E.2.2 Convergence of Error Terms
We have just seen in Lemma E.25 all individually have expected value $G(w)$. To establish the result we will verify Lyapounov’s condition, that is, we need to check that, for some $\delta>0$,
@@
We will use this version of the central limit theorem to establish uniform convergence properties of the random variables $\psi_{n}(w)$ over compact subsets of $\accentset{\circ}{W}\ \backslash\ W_{0}$. Our goal is an analogue of [Watanabe, 2009, Theorem 5.9]. More specifically we will prove the following.
Theorem E.28.
@@
Our proof strategy – following Watanabe – will combine Prokhorov’s theorem, telling us that is we can prove $\psi_{n}|_{K}$ is a uniformly tight sequence then it admits a convergent subsequence, with Proposition E.27 telling us that the sequence $\psi_{n}(w)$ converges pointwise for $w\in\accentset{\circ}{W}\ \backslash\ W_{0}$. We will begin with a sequence of lemmas that will together prove uniform tightness. Throughout these lemmas we fix a compact subset $K\subseteq\accentset{\circ}{W}\ \backslash\ W_{0}$.
Lemma E.29.
@@
We will generalize the argument given by in [Watanabe, 2009, Theorem 5.7, 5.8]. Recall that
@@
Let us consider any analytic function $a\colon\mathbb{R}\to\mathbb{R}$ and consider the argument in [Watanabe, 2009, Theorem 5.7]. Without the assumption that the random variables $X_{i}$ are identically distributed Watanabe’s argument still establishes
@@
We will prove this in two steps. First choose an even integer $s>2$ as Lemma E.30 and use the lemma to find $C>0$ so that
@@
Remark E.32.
The argument given for this claim in [Watanabe, 2009, Example 5.3] requires an additional step since closed balls in the $L^{\infty}$ topology on infinite spaces are not compact. This point and the correction used above were communicated to us by Rohan Hitchcock.
Corollary E.33.
@@
Proof of Theorem E.28.
Applying Prokhorov’s theorem, Corollary E.33 tells us that $\psi_{n}(w)$ has a convergent subsequence. Now Proposition E.27 says that $\psi_{n}(w)$ converges in distribution pointwise for $w\in K$, and therefore $\psi_{n}(w)$ converges in $C^{\omega}(K)$ to a random variable $\psi|_{K}$. ∎
E.2.3 Resolution of Singularities
Our next goal is to establish part (1) of Theorem E.24. That is, we will give a local (in parameter space) description of the importance weighted estimator $G_{n}$ and the empirical process $\psi_{n}$ in a standard form that is comparatively easy to manipulate. We will do this by fixing a log resolution of the zero set $W_{0}\subseteq W$ of the loss function. We recall the fundamental construction as originally developed by Hironaka Hironaka [1964].
Definition E.34.
@@ E.2.3 Resolution of Singularities
Choose a point $w^{*}\in W_{0}$, a local log resolution $\varpi$ and a chart $U$ around this minimum. Let $U_{0}=\varpi^{-1}(W_{0})\cap U$ and let $K$ be a compact subset of $\accentset{\circ}{U}\ \backslash\ U_{0}$. In this chart our importance weighted loss may be usefully expressed in a standard way – in normal crossings form – parallelling [Watanabe, 2009, Main Theorem 1]. We will begin by representing the random variable $f$ from Definition E.18 in normal crossings form.
Proposition E.35.
@@
We obtain $a,b$ by forming the Taylor expansion in $u$ of the analytic function $f(\varpi(u),w^{\prime},x)$ and splitting the resulting sum into the sum of those Taylor terms of order at least $u_{1}^{\lceil\ell/2\rceil}u_{2}^{k_{2}}\cdots u_{d}^{k_{d}}$, and the sum of those Taylor terms of smaller degree in at least one variable. Having done so, by Lemma E.25 and the definition of the log resolution we have
@@
By Theorem E.28 and Proposition E.27 we know that $\sqrt{n}(G(w)-G_{n}(w))\to\mathcal{N}(0,\sigma(w,w^{*}))$ in distribution uniformly over compact subspaces of $\accentset{\circ}{W}\ \backslash\ W_{0}$. In particular for the resolution $\varpi$ we have
@@
E.2.4 Free Energy Asymptotics
Let us now return to Theorem E.24, in this section we will address part (2), concerning the asymptotic behavior of the generalized posterior distribution. In order to establish clear asymptotics we will need to guarantee that the lower order terms in the importance weighted loss estimators – the terms $\omega_{n}$ from Theorem E.36 – do not contribute to the posterior. This is not likely to hold in general, we will establish it under the condition that Assumption E.16 holds, i.e. that if $w^{*}\in W$ minimizes the loss then the set
@@ E.2.4 Free Energy Asymptotics
Proposition E.37.
If $q_{w}$ is a probability distribution for all $w$ and Assumption E.16 holds then $\omega_{n}=0$ for all $n$.
Remark E.38.
Proposition E.37 is very close to the condition of relatively finite variance that appears in Watanabe [2018]. In the general case where $q_{w}$ is merely a signed measure (which includes the density estimation problem originally considered by Watanabe) the relatively finite variance condition is what is needed to establish $\omega_{n}=0$. This is, however, implied by Assumption E.16 in the positive case. Indeed, recall that the function $f$ has relatively finite variance if
@@
for all $w^{\prime}\in\accentset{\circ}{W}\ \backslash\ W_{0}$. If we work in a local chart $U$ in a resolution around a minimum $w^{*}$ of the loss then $\mathbb{E}(f(w,w^{\prime},x))=u^{2k}$, and we will see in the proof of Proposition E.37 that under the assumption $f(w,w^{\prime},x)$ also has Taylor expansion with leading term of order $2k$, so the ratio of the variance and expectation also converges to zero along as $u\to 0$.
We will assume below that $q_{w}$ is a probability distribution for all $w$; in particular that it is non-negative. If $q_{w}=-\log p_{w}$ for a model $p_{w}$, as in the distribution learning example E.19, this is not generally true. In this case Proposition E.37 still holds in the realizable case by [Watanabe, 2009, Theorem 6.1] or under a more general relative finite variance condition as in [Watanabe, 2018, Theorem 8].
Lemma E.39.
Let $w^{*}\in W$ minimize the loss. If $q_{w}$ is a probability distribution for all $w$ and Assumption E.16 holds then
@@
Since $h(x)\geq 0$ and $q_{w^{*}}(x)>0$ almost surely, if $g(x)q_{w^{*}}(x)-h(x)\geq 0$ then $g(x)\geq 0$. If $g(x)=0$ then $g(x)q_{w^{*}}(x)-h(x)<0$, so by Assumption E.16 we conclude that $g(x)>0$ with probability one. ∎
Proof of Proposition E.37.
@@
where we use absolute convergence of the Taylor series to exchange the sum and the integral. By Lemma E.39 $g(x)>0$ on $X_{w^{*}}$ almost everywhere. Since $b_{\alpha}(x)$ are the Taylor coefficients of a non-negative function, non-zero for some $x\in X_{w^{*}}$, the leading Taylor coefficient must be positive, and therefore $b_{\alpha}=0$ if $0<\alpha<2k$. The constant term $b_{0}(x)=q_{w^{*}}(x)$. Therefore
@@
Finally, Assumption E.16 implies that $f(w^{*},w^{\prime},x)=0$, and therefore the Taylor series of $f(\varpi(u),w^{\prime},x)$ in $u$ is concentrated in degrees at least $2k$, and therefore in particular $\omega_{n}=0$ for all $n$. ∎
@@
Theorem E.40.
Suppose that $q_{w}$ is a family of probability distributions and Assumption E.16 holds. Let $U\subseteq W$ be an open subset, and let $(\lambda,m)=(\mathrm{rlct}_{U}(G),\mathrm{rlcm}_{U}(G))$ denote the real log canonical threshold and multiplicity of the loss function $G$ on the subset $U$. Then
@@
Remark E.41.
One can give an explicit integral formula for $C(\beta,\varphi)$ as a sum over local charts in a log resolution, see [Watanabe, 2018, Section 6.3] or [Watanabe, 2009, Section 6.2] for such an expression.
@@
Given Proposition E.37 and Theorem E.36 this follows from Watanabe’s argument from [Watanabe, 2009, Theorem 6.7]. ∎
Corollary E.42.
@@
E.2.5 Expected Total Loss and the Widely Applicable Bayesian Information Criterion
We will now conclude our technical arguments by establishing part (3) of Theorem E.24, the analogue of the widely applicable Bayesian information criterion Watanabe [2013], explaining how the generalized posterior selects between different critical regions of $W_{0}$ as the number $n$ of datapoints increases.
@@ E.2.5 Expected Total Loss and the Widely Applicable Bayesian Information Criterion
Theorem E.43.
Suppose that $q_{w}$ is a family of probability distributions and Assumption E.16 holds. Let $\beta\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a positive function on the natural numbers such that $\beta(n)$ converges as $n\to\infty$. Let $w^{*}$ be a local minimum of the loss and let $U$ be an open neighborhood of $w^{*}$ in $W$. Then
@@
We will follow the argument in [Watanabe, 2013, Theorem 4] closely. It suffices to work locally in $W$; choose a local log resolution $\varpi\colon\mathcal{M}\to W$ and a local chart $V$ with local coordinate $u$ centered around a preimage of $w^{*}$. Choose the chart so that $G_{0}=G(\varpi(0))$ is the minimum value of $G\circ\varpi$ on $V$. Let $\widetilde{f}(w,w^{\prime},x)=f(w,w^{\prime},x)-G_{0}$ and $\widetilde{G}_{n}=G_{n}=G_{0}$. Then
@@
We will study the numerator and denominator of our WBIC expression separately. Since we have established Proposition E.37 and Theorem E.36 Watanabe’s calculations of these two expressions continue to apply in our setting; that is, by the argument in the proof of [Watanabe, 2013, Theorem 4] we have asymptotic behavior
@@
In Watanabe [2013] Watanabe uses the specific form $\beta=\beta_{0}/\log(n)$ for comparison with the free energy of the Bayesian posterior. That is for $\mathcal{F}_{n,\beta_{0}}(U)$ as defined in Corollary E.42 we have
by Theorem E.43. The term widely applicable Bayesian information criterion (WBIC) is used for this expression since it generalizes the classical Bayesian information criterion Schwarz [1978], which this result recovers in the regular case.
Appendix F Theory of LLC Estimation in Reinforcement Learning
Definition F.1.
Remark F.2.
@@ Appendix F Theory of LLC Estimation in Reinforcement Learning
As we have seen from Theorem E.40 the local learning coefficient controls the local behaviour of the generalized Bayesian posterior distribution near the point $w^{*}$ in the large $n$ limit. In general when $\dim W$ is very large we cannot expect to compute $\lambda$ exactly, so in this section we will discuss statistical estimators that we will use to substitute for it. Our estimators are based on the local learning coefficient estimators developed in Lau et al. [2025].
F.1 Asymptotically Unbiased Estimators
We will begin with those estimators for which we can establish good theoretical properties rigorously, and generalize from there to estimators over which we have less theoretical control but which are more practical to estimate in realistic examples. To start with we can use the WBIC – Theorem E.43 – to define asymptotically unbiased estimators for the local learning coefficient.
Given a sequence of samples $x_{i}\sim q_{w_{i}}$ as in Definition E.18, choose $\beta>0$ and let $\Omega_{n,\beta}$ be the normalized posterior distribution from Definition 2.1. Choose an open neighborhood $U$ of $w^{*}$ and a natural number $T$ and let $y_{1},\ldots,y_{T}$ be independent random variables from the restriction of the generalized tempered posterior to $U$: $y_{j}\sim\Omega_{n,\beta}|_{U}$.
Definition F.3.

### v2

@@ D.2.2 Convergence of Error Terms
We have just seen in Lemma D.25 all individually have expected value $G(w)$. To establish the result we will verify Lyapounov’s condition, that is, we need to check that, for some $\delta>0$,
@@
We will use this version of the central limit theorem to establish uniform convergence properties of the random variables $\psi_{n}(w)$ over compact subsets of $\accentset{\circ}{W}\ \backslash\ W_{0}$. Our goal is an analogue of Watanabe [2009, Theorem 5.9]. More specifically we will prove the following.
Theorem D.28.
@@
Our proof strategy – following Watanabe – will combine Prokhorov’s theorem, telling us that is we can prove $\psi_{n}|_{K}$ is a uniformly tight sequence then it admits a convergent subsequence, with Proposition D.27 telling us that the sequence $\psi_{n}(w)$ converges pointwise for $w\in\accentset{\circ}{W}\ \backslash\ W_{0}$. We will begin with a sequence of lemmas that will together prove uniform tightness. Throughout these lemmas we fix a compact subset $K\subseteq\accentset{\circ}{W}\ \backslash\ W_{0}$.
Lemma D.29.
@@
We will generalize the argument given by in Watanabe [2009, Theorems 5.7, 5.8]. Recall that
@@
Let us consider any analytic function $a\colon\mathbb{R}\to\mathbb{R}$ and consider the argument in Watanabe [2009, Theorem 5.7]. Without the assumption that the random variables $X_{i}$ are identically distributed Watanabe’s argument still establishes
@@
We will prove this in two steps. First choose an even integer $s>2$ as Lemma D.30 and use the lemma to find $C>0$ so that
@@
Remark D.32.
The argument given for this claim in Watanabe [2009, Example 5.3] requires an additional step since closed balls in the $L^{\infty}$ topology on infinite spaces are not compact. This point and the correction used above were communicated to us by Rohan Hitchcock.
Corollary D.33.
@@
Proof of Theorem D.28.
Applying Prokhorov’s theorem, Corollary D.33 tells us that $\psi_{n}(w)$ has a convergent subsequence. Now Proposition D.27 says that $\psi_{n}(w)$ converges in distribution pointwise for $w\in K$, and therefore $\psi_{n}(w)$ converges in $C^{\omega}(K)$ to a random variable $\psi|_{K}$. ∎
D.2.3 Resolution of Singularities
Our next goal is to establish part (1) of Theorem D.24. That is, we will give a local (in parameter space) description of the importance weighted estimator $G_{n}$ and the empirical process $\psi_{n}$ in a standard form that is comparatively easy to manipulate. We will do this by fixing a log resolution of the zero set $W_{0}\subseteq W$ of the loss function. We recall the fundamental construction as originally developed by Hironaka [1964].
Definition D.34.
@@ D.2.3 Resolution of Singularities
Choose a point $w^{*}\in W_{0}$, a local log resolution $\varpi$ and a chart $U$ around this minimum. Let $U_{0}=\varpi^{-1}(W_{0})\cap U$ and let $K$ be a compact subset of $\accentset{\circ}{U}\ \backslash\ U_{0}$. In this chart our importance weighted loss may be usefully expressed in a standard way – in normal crossings form – parallelling Watanabe [2009, Main Theorem 1]. We will begin by representing the random variable $f$ from Definition D.18 in normal crossings form.
Proposition D.35.
@@
We obtain $a,b$ by forming the Taylor expansion in $u$ of the analytic function $f(\varpi(u),w^{\prime},x)$ and splitting the resulting sum into the sum of those Taylor terms of order at least $u_{1}^{\lceil\ell/2\rceil}u_{2}^{k_{2}}\cdots u_{d}^{k_{d}}$, and the sum of those Taylor terms of smaller degree in at least one variable. Having done so, by Lemma D.25 and the definition of the log resolution we have
@@
By Theorem D.28 and Proposition D.27 we know that $\sqrt{n}(G(w)-G_{n}(w))\to\mathcal{N}(0,\sigma(w,w^{*}))$ in distribution uniformly over compact subspaces of $\accentset{\circ}{W}\ \backslash\ W_{0}$. In particular for the resolution $\varpi$ we have
@@
D.2.4 Free Energy Asymptotics
Let us now return to Theorem D.24, in this section we will address part (2), concerning the asymptotic behavior of the generalized posterior distribution. In order to establish clear asymptotics we will need to guarantee that the lower order terms in the importance weighted loss estimators – the terms $\omega_{n}$ from Theorem D.36 – do not contribute to the posterior. This is not likely to hold in general, we will establish it under the condition that Assumption D.16 holds, i.e. that if $w^{*}\in W$ minimizes the loss then the set
@@ D.2.4 Free Energy Asymptotics
Proposition D.37.
If $q_{w}$ is a probability distribution for all $w$ and Assumption D.16 holds then $\omega_{n}=0$ for all $n$.
Remark D.38.
Proposition D.37 is very close to the condition of relatively finite variance that appears in Watanabe [2018]. In the general case where $q_{w}$ is merely a signed measure (which includes the density estimation problem originally considered by Watanabe) the relatively finite variance condition is what is needed to establish $\omega_{n}=0$. This is, however, implied by Assumption D.16 in the positive case. Indeed, recall that the function $f$ has relatively finite variance if
@@
for all $w^{\prime}\in\accentset{\circ}{W}\ \backslash\ W_{0}$. If we work in a local chart $U$ in a resolution around a minimum $w^{*}$ of the loss then $\mathbb{E}(f(w,w^{\prime},x))=u^{2k}$, and we will see in the proof of Proposition D.37 that under the assumption $f(w,w^{\prime},x)$ also has Taylor expansion with leading term of order $2k$, so the ratio of the variance and expectation also converges to zero along as $u\to 0$.
We will assume below that $q_{w}$ is a probability distribution for all $w$; in particular that it is non-negative. If $q_{w}=-\log p_{w}$ for a model $p_{w}$, as in the distribution learning example D.19, this is not generally true. In this case Proposition D.37 still holds in the realizable case by Watanabe [2009, Theorem 6.1] or under a more general relative finite variance condition as in Watanabe [2018, Theorem 8].
Lemma D.39.
Let $w^{*}\in W$ minimize the loss. If $q_{w}$ is a probability distribution for all $w$ and Assumption D.16 holds then
@@
Since $h(x)\geq 0$ and $q_{w^{*}}(x)>0$ almost surely, if $g(x)q_{w^{*}}(x)-h(x)\geq 0$ then $g(x)\geq 0$. If $g(x)=0$ then $g(x)q_{w^{*}}(x)-h(x)<0$, so by Assumption D.16 we conclude that $g(x)>0$ with probability one. ∎
Proof of Proposition D.37.
@@
where we use absolute convergence of the Taylor series to exchange the sum and the integral. By Lemma D.39 $g(x)>0$ on $X_{w^{*}}$ almost everywhere. Since $b_{\alpha}(x)$ are the Taylor coefficients of a non-negative function, non-zero for some $x\in X_{w^{*}}$, the leading Taylor coefficient must be positive, and therefore $b_{\alpha}=0$ if $0<\alpha<2k$. The constant term $b_{0}(x)=q_{w^{*}}(x)$. Therefore
@@
Finally, Assumption D.16 implies that $f(w^{*},w^{\prime},x)=0$, and therefore the Taylor series of $f(\varpi(u),w^{\prime},x)$ in $u$ is concentrated in degrees at least $2k$, and therefore in particular $\omega_{n}=0$ for all $n$. ∎
@@
Theorem D.40.
Suppose that $q_{w}$ is a family of probability distributions and Assumption D.16 holds. Let $U\subseteq W$ be an open subset, and let $(\lambda,m)=(\mathrm{rlct}_{U}(G),\mathrm{rlcm}_{U}(G))$ denote the real log canonical threshold and multiplicity of the loss function $G$ on the subset $U$. Then
@@
Remark D.41.
One can give an explicit integral formula for $C(\beta,\varphi)$ as a sum over local charts in a log resolution, see Watanabe [2018, Section 6.3] or Watanabe [2009, Section 6.2] for such an expression.
@@
Given Proposition D.37 and Theorem D.36 this follows from Watanabe’s argument from Watanabe [2009, Theorem 6.7]. ∎
Corollary D.42.
@@
D.2.5 Expected Total Loss and the Widely Applicable Bayesian Information Criterion
We will now conclude our technical arguments by establishing part (3) of Theorem D.24, the analogue of the widely applicable Bayesian information criterion Watanabe [2013], explaining how the generalized posterior selects between different critical regions of $W_{0}$ as the number $n$ of datapoints increases.
@@ D.2.5 Expected Total Loss and the Widely Applicable Bayesian Information Criterion
Theorem D.43.
Suppose that $q_{w}$ is a family of probability distributions and Assumption D.16 holds. Let $\beta\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a positive function on the natural numbers such that $\beta(n)$ converges as $n\to\infty$. Let $w^{*}$ be a local minimum of the loss and let $U$ be an open neighborhood of $w^{*}$ in $W$. Then
@@
We will follow the argument in Watanabe [2013, Theorem 4] closely. It suffices to work locally in $W$; choose a local log resolution $\varpi\colon\mathcal{M}\to W$ and a local chart $V$ with local coordinate $u$ centered around a preimage of $w^{*}$. Choose the chart so that $G_{0}=G(\varpi(0))$ is the minimum value of $G\circ\varpi$ on $V$. Let $\widetilde{f}(w,w^{\prime},x)=f(w,w^{\prime},x)-G_{0}$ and $\widetilde{G}_{n}=G_{n}=G_{0}$. Then
@@
We will study the numerator and denominator of our WBIC expression separately. Since we have established Proposition D.37 and Theorem D.36 Watanabe’s calculations of these two expressions continue to apply in our setting; that is, by the argument in the proof of Watanabe [2013, Theorem 4] we have asymptotic behavior
@@
Remark D.44.
Watanabe [2013] uses the specific form $\beta=\beta_{0}/\log(n)$ for comparison with the free energy of the Bayesian posterior. That is for $\mathcal{F}_{n,\beta_{0}}(U)$ as defined in Corollary D.42 we have
by Theorem D.43. The term widely applicable Bayesian information criterion (WBIC) is used for this expression since it generalizes the classical Bayesian information criterion [Schwarz, 1978], which this result recovers in the regular case.
Appendix E Theory of LLC Estimation in Reinforcement Learning
In this appendix, we offer a self-contained introduction to LLC estimation, following Lau et al. [2025], but adapted to the RL context.
Definition E.1.
@@ Appendix E Theory of LLC Estimation in Reinforcement Learning
As we have seen from Theorem D.40 the local learning coefficient controls the local behaviour of the generalized Bayesian posterior distribution near the point $w^{*}$ in the large $n$ limit. In general when $\dim W$ is very large we cannot expect to compute $\lambda$ exactly, so in this section we will discuss statistical estimators that we will use to substitute for it. Our estimators are based on the local learning coefficient estimators developed in Lau et al. [2025].
E.1 Asymptotically Unbiased Estimators
We will begin with those estimators for which we can establish good theoretical properties rigorously, and generalize from there to estimators over which we have less theoretical control but which are more practical to estimate in realistic examples. To start with we can use the WBIC – Theorem D.43 – to define asymptotically unbiased estimators for the local learning coefficient.
Given a sequence of samples $x_{i}\sim q_{w_{i}}$ as in Definition D.18, choose $\beta>0$ and let $\Omega_{n,\beta}$ be the normalized posterior distribution from Definition 3.1. Choose an open neighborhood $U$ of $w^{*}$ and a natural number $T$ and let $y_{1},\ldots,y_{T}$ be independent random variables from the restriction of the generalized tempered posterior to $U$: $y_{j}\sim\Omega_{n,\beta}|_{U}$.
Definition E.3.